Problem 14. A 7-kg and a 11-kg box are suspended by a large pulley that has a mass of 3 kg and a radius of 0.15 m. The masses start 3.5 m apart, and the system is released from rest. What is the speed of the boxes when they pass one another? 11kg: 3.5m 7kg
Answer #1

3.5m Let the acceleration of system is a. and Tension in string is T, and T. Hence, appaying Newtons law, T-79=79 0 g-T=hla10 Hence the speed of bobes when they pass one another applying equation of motion. Taval 12=42 +298 - v2 = 02 + 2x & g x 35